# Enzyme-Substrate Complex Formation Modulates Diffusion-Driven Patterning in Metabolic Pathways

### v1

@@ 3.2. Equilibrium Analysis and Derivation of the Non-degenerated Model
To obtain the equilibrium concentrations $C_{0}^{*}$, $C_{1}^{*}$, and $C_{E_{1}S_{0}}^{*}$, we first note that from the first equation of system (1.2) one can express $C_{0}=C_{0}^{*}-\dfrac{k_{d_{1}}n_{1}\,C_{E_{1}S_{0}}}{\alpha_{0}}.$ From the second equation, under the simplifying assumption $k_{a_{1}}n_{1}=k_{d_{1}}+k_{cat}$, the model becomes degenerate. Solving the third equation for $C_{1}$ gives $C_{1}=\dfrac{k_{cat}}{k_{2}n_{2}+\beta}\,C_{E_{1}S_{0}}.$
Since the system is thus degenerate, to obtain a reduced model with independent variables we employ the quasi–steady–state approximation (QSSA). This assumption states that the enzyme–substrate complex $C_{E_{1}S_{0}}$ relaxes much faster than the concentrations of $C_{0}$ and $C_{1}$, i.e. $\dfrac{\partial C_{E_{1}S_{0}}}{\partial t}\approx 0.$ Hence, from the second equation of system (1.2) we can write $C_{E_{1}S_{0}}=\dfrac{k_{a_{1}}n_{1}C_{0}+D\nabla^{2}C_{E_{1}S_{0}}}{k_{d_{1}}+k_{cat}}.$
Assuming that the diffusive contribution $D\nabla^{2}C_{E_{1}S_{0}}$ is negligible compared with the reaction term $k_{a_{1}}n_{1}C_{0}$, we obtain the approximation
$C_{E_{1}S_{0}}\approx\dfrac{k_{a_{1}}n_{1}}{k_{d_{1}}+k_{cat}}\,C_{0}.$ (3.2)
Substituting (3.2) into the remaining equations of system (1.2) yields the following reduced or non-degenerated model:
@@
The non-degenerated system (3.3) retains the same structure as the primary model (1.1), but with modified effective coefficients that encapsulate the influence of the enzyme–substrate complex through the QSSA. The variable $C_{E_{1}S_{0}}$ has been successfully eliminated, and the resulting model now involves only the substrate and intermediate/product concentrations $(C_{0},C_{1})$. This formulation allows for direct comparison with the simplified model (1.1), while maintaining the mechanistic effects of enzyme–substrate binding and catalytic conversion.
3.3. Linear Analysis of enzyme-substrate complex
The model (3.3) has an unique coexistence steady state as:
$u_{0}^{*}=\dfrac{\alpha_{0}u_{in}^{*}(k_{d_{1}}+k_{cat})}{\alpha_{0}(k_{d_{1}}+k_{cat})+k_{a_{1}}k_{d_{1}}n_{1}^{2}},\;\;u_{1}^{*}=\dfrac{k_{a_{1}}k_{cat}n_{1}\alpha_{0}u_{in}^{*}}{(k_{d_{1}}+k_{cat})(k_{a_{1}}k_{d_{1}}n_{1}^{2})}$ (3.4)
@@ 3.3. Linear Analysis of enzyme-substrate complex
$J(E^{*})=\begin{pmatrix}-\alpha_{0}-\dfrac{k_{a_{1}}k_{d_{1}}n^{2}}{k_{d_{1}}+k_{cat}}&&0\\ \dfrac{k_{a_{1}}k_{cat}n_{1}}{k_{d_{1}}+k_{cat}}&&-\beta-k_{2}n_{2}\end{pmatrix},\;\;\;\mathcal{D}(E^{*})=\begin{pmatrix}\gamma_{0}&&-\gamma u_{0}^{*}\\ 0&&\gamma_{1}\end{pmatrix}$ (3.5)
@@
A qualitative comparison between the kinetic dynamics of the simplified model (1.1) and the reduced non-degenerated model (3.3) is shown in Figure 1. The phase–plane analysis highlights several key differences: (i) the nullclines of the non-degenerated model intersect at a steady state with higher intermediate concentration, (ii) trajectories relax more slowly toward equilibrium, reflecting the additional nonlinear feedback introduced by enzyme–substrate binding, and (iii) the vector fields of the two systems display distinct curvature, indicating that the mechanistic correction introduced by the QSSA reduction significantly alters the underlying kinetics.
@@ 4. Turing region
To understand how parameters influence on appearance of Turing instability of both models, we need to obtain Turing regions. In order to achieve this, we need to consider positivity conditions of both $E^{*}$, stability conditions in absence of diffusion. According to the (2.1) and (3.4) coexistence steady states are always positive for choosing any positive values of parameters. Moreover, it is locally stable for choosing any positive values of parameters. Thus, there are only Turing conditions (2.10), and (3.6). According to these conditions, we investigate variation of parameters in plane $(n_{1},d),\;(n_{1},\gamma)$. Figure 2, demonstrates that for the given data set, the bifurcation parameters of models (1.1) and (3.3) are $d_{c}=2.7190$ and $\gamma_{c}=5.4035$ correspondingly.
@@ 5. Weakly nonlinear Analysis of models
In this section we perform a weakly nonlinear analysis for the model (1.1) to predict the amplitude of the pattern near the Turing threshold. The weakly nonlinear analysis is based on the multiple scale methods [12]. The similar strategy has applied on the model (3.3). Since near to the bifurcation the amplitude of the pattern (diffusion-driven instability) has slow temporal scale, then a new temporal scale is defined.
@@ 7. Conclusion
This work examined how enzymatic reaction kinetics and diffusion-driven transport interact to generate spatial metabolic organization. The linear stability and Turing analyses showed that diffusion and curvature-driven cross-diffusion alone can destabilize otherwise stable homogeneous steady states, producing spatial patterns through local enzyme–substrate feedback. Incorporating explicit complex formation introduces additional kinetic structure that modifies both the reaction balance and the effective diffusion terms, thereby shifting the Turing threshold and altering the size and location of the instability region. The quasi-steady-state reduction yields a non-degenerate model that preserves the essential nonlinear coupling between catalysis and compartmentalization while enabling a complete weakly nonlinear (WNL) analysis.
Theoretical predictions from the WNL framework were strongly supported by numerical simulations, which showed excellent agreement in both pattern amplitude and spatial profile. The comparison between the simplified and non-degenerate models revealed systematic kinetic and spatial differences: explicit enzyme–substrate binding shifts the equilibrium toward higher intermediate concentrations, slows relaxation toward steady state, and produces spatial patterns of smaller amplitude. These results highlight the buffering role of reversible complex formation and demonstrate how enzymatic binding kinetics regulate the strength of the feedback loop that drives Turing instability.
Beyond their mathematical implications, these findings have a natural interpretation in the context of liquid–liquid phase separation (LLPS). Inside biomolecular condensates, enzymes and metabolites experience altered mobility, molecular crowding, and transient binding events that modulate both catalytic flux and spatial organization. The simplified model corresponds to a regime of strong local co-localization where binding is effectively instantaneous, whereas the non-degenerate model captures a more realistic buffered regime in which reversible complex formation attenuates spatial feedback. The resulting differences in pattern amplitude and stability mirror known biophysical roles of LLPS in regulating metabolite gradients, enhancing local reaction efficiency, and preventing uncontrolled accumulation of intermediates.
Overall, this study establishes a quantitative link between enzymatic binding kinetics, diffusion-driven instabilities, and mesoscale spatial organization such as LLPS. The framework developed here can be readily extended to more complex metabolic pathways, multi-enzyme assemblies, and other compartmentalized biochemical systems, offering a versatile foundation for understanding how molecular self-organization shapes cellular metabolism and morphogenesis.

### v2

@@ 3.2. Equilibrium Analysis and Derivation of the Non-degenerated Model
To obtain the equilibrium concentrations $C_{0}^{*}$, $C_{1}^{*}$, and $C_{E_{1}S_{0}}^{*}$, we first rewrite the steady-state conditions corresponding to system (1.2). From the first equation we obtain the linear relation $C_{0}=C_{0}^{*}-\frac{k_{d_{1}}n_{1}}{\alpha_{0}}\,C_{E_{1}S_{0}}.$ Under the simplifying condition $k_{a_{1}}n_{1}=k_{d_{1}}+k_{cat}$, the second equation becomes singular and the homogeneous kinetic system is degenerate along a one-dimensional manifold of equilibria. Solving the third equation for the intermediate $C_{1}$ yields $C_{1}=\frac{k_{cat}}{k_{2}n_{2}+\beta}\,C_{E_{1}S_{0}}.$
Because of this degeneracy, it is useful to construct a reduced model in which the enzyme–substrate complex is expressed in terms of the substrate concentration. To this end, we adopt a quasi-equilibrium approximation for $C_{E_{1}S_{0}}$ by assuming that, on the spatial and temporal scales of interest, its diffusive contribution is small compared to the local reaction term. Neglecting the term $D\nabla^{2}C_{E_{1}S_{0}}$ in the second equation of (1.2) then gives the algebraic approximation $(k_{d_{1}}+k_{cat})\,C_{E_{1}S_{0}}\approx k_{a_{1}}n_{1}\,C_{0},$ and thus
$C_{E_{1}S_{0}}\approx\frac{k_{a_{1}}n_{1}}{k_{d_{1}}+k_{cat}}\,C_{0}.$ (3.2)
Substituting (3.2) into the remaining equations of system (1.2) yields the following reduced, or non-degenerated, two-variable model:
@@
The non-degenerated system (3.3) has the same structure as the primary model (1.1), but with modified effective coefficients that incorporate the contribution of the enzyme–substrate complex via the quasi-equilibrium approximation (3.2). In this formulation, $C_{E_{1}S_{0}}$ no longer appears as an independent variable, while the mechanistic effects of binding and catalytic conversion are retained through the renormalized reaction terms. This reduced model provides a convenient comparison to the primary pathway without altering the full three-species dynamics used for the pattern-formation analysis.
3.3. Linear Analysis of the Non-degenerated Reduced Model
The reduced non-degenerated system (3.3), obtained via the quasi-equilibrium approximation of the enzyme-substrate complex, admits a unique positive coexistence steady state given by
$u_{0}^{*}=\dfrac{\alpha_{0}u_{in}^{*}(k_{d_{1}}+k_{cat})}{\alpha_{0}(k_{d_{1}}+k_{cat})+k_{a_{1}}k_{d_{1}}n_{1}^{2}},\qquad u_{1}^{*}=\dfrac{k_{a_{1}}k_{cat}n_{1}\alpha_{0}u_{in}^{*}}{(k_{d_{1}}+k_{cat})(k_{a_{1}}k_{d_{1}}n_{1}^{2})}.$ (3.4)
All parameters are assumed to be positive.
@@ 3.3. Linear Analysis of the Non-degenerated Reduced Model
$J(E^{*})=\begin{pmatrix}-\alpha_{0}-\dfrac{k_{a_{1}}k_{d_{1}}n^{2}}{k_{d_{1}}+k_{cat}}&&0\\ \dfrac{k_{a_{1}}k_{cat}n_{1}}{k_{d_{1}}+k_{cat}}&&-\beta-k_{2}n_{2}\end{pmatrix},\;\;\;\mathcal{D}(E^{*})=\begin{pmatrix}\gamma_{0}&&-\gamma u_{0}^{*}\\ 0&&\gamma_{1}\end{pmatrix}$. (3.5)
@@
A qualitative comparison between the kinetic dynamics of the simplified model (1.1) and the reduced non-degenerated model (3.3) is shown in Figure 1. The phase–plane analysis highlights several key differences: (i) the nullclines of the non-degenerated model intersect at a steady state with higher intermediate concentration, (ii) trajectories relax more slowly toward equilibrium, reflecting the additional nonlinear feedback introduced by enzyme–substrate binding, and (iii) the vector fields of the two systems display distinct curvature, indicating that the mechanistic correction introduced by the approximation reduction significantly alters the underlying kinetics.
@@ 4. Turing region
To understand how parameters influence the emergence of Turing instability in the full enzymatic model, and to compare these effects with those predicted by the reduced non-degenerated system, we investigate the corresponding Turing regions in parameter space. In order to achieve this, we need to consider positivity conditions of both $E^{*}$, stability conditions in absence of diffusion. According to the (2.1) and (3.4) for the parameter ranges considered in this study, the coexistence steady states of both models are positive and locally asymptotically stable in the absence of diffusion. Thus, there are only Turing conditions (2.10), and (3.6). According to these conditions, we investigate variation of parameters in plane $(n_{1},d),\;(n_{1},\gamma)$. Figure 2, demonstrates that for the given data set, the bifurcation parameters of models (1.1) and (3.3) are $d_{c}=2.7190$ and $\gamma_{c}=5.4035$ correspondingly.
@@ 5. Weakly nonlinear Analysis of models
In this section we perform a weakly nonlinear analysis for the model (1.1) to predict the amplitude of the pattern near the Turing threshold. The weakly nonlinear analysis is based on the multiple scale methods (Gambino et al. (2012)). The similar strategy has applied on the model (3.3). Since near to the bifurcation the amplitude of the pattern (diffusion-driven instability) has slow temporal scale, then a new temporal scale is defined.
@@ 7. Conclusion
This work examined how enzymatic reaction kinetics and diffusion-driven transport interact to generate spatial metabolic organization. Linear stability and Turing analyses showed that diffusion and curvature-driven cross-diffusion can destabilize otherwise stable homogeneous steady states, producing spatial patterns through local enzyme–substrate feedback. Incorporating explicit enzyme–substrate complex formation introduces additional kinetic structure that modifies reaction balance and effective transport, thereby shifting the Turing threshold and altering the size and location of the instability region. The resulting reduced non-degenerate formulation removes the structural degeneracy associated with a manifold of equilibria and enables a complete weakly nonlinear (WNL) analysis while preserving the essential nonlinear coupling between catalysis and spatial organization.
Theoretical predictions from the WNL framework are strongly supported by numerical simulations, which show excellent agreement in both pattern amplitude and spatial profile. Comparison between the simplified and non-degenerate models reveals systematic kinetic and spatial differences: explicit enzyme–substrate binding shifts the homogeneous equilibrium toward higher intermediate concentrations, slows relaxation toward steady state, and leads to spatial patterns of reduced amplitude. These results highlight the buffering role of reversible complex formation and demonstrate how enzymatic binding kinetics regulate the strength of the feedback loop driving diffusion-driven instability.
Beyond their mathematical implications, these findings admit a natural interpretation in the context of liquid–liquid phase separation (LLPS). Within biomolecular condensates, enzymes and metabolites experience altered mobility, molecular crowding, and transient binding interactions that jointly modulate catalytic flux and spatial organization. The simplified model captures an effective, coarse-grained description of enzymatic feedback, whereas the non-degenerate formulation reflects a buffered regime in which reversible complex formation attenuates spatial feedback. The resulting differences in pattern amplitude and stability mirror known biophysical roles of LLPS in regulating metabolite gradients, enhancing reaction efficiency, and preventing uncontrolled accumulation of intermediates.
Overall, this study establishes a quantitative link between enzymatic binding kinetics, diffusion-driven instabilities, and mesoscale spatial organization such as LLPS. The framework developed here can be extended to more complex metabolic pathways, multi-enzyme assemblies, and other compartmentalized biochemical systems, providing a versatile foundation for understanding how molecular self-organization shapes cellular metabolism and morphogenesis.
The degenerate structure of the full enzymatic model suggests rich reduced dynamics near criticality. Exploring rigorous invariant-manifold reductions and more detailed enzyme–substrate formulations is an exciting direction that we are actively pursuing.
@@ 8. Acknowledgments
Statements and Declarations
Competing interests: The author declares that there are no competing interests.
Data availability: No new experimental data were generated in this study. All simulation codes and any numerical data supporting the findings are available from the corresponding author on reasonable request.
Funding: No external funding was received for this research.
Ethics approval and consent to participate: Not applicable.
Authors’ contributions: The author developed the mathematical model, performed all analyses and simulations, and wrote the entire manuscript.